40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 16228     Entry time: Tue Jun 29 17:42:06 2021
Author: Anchal, Paco, Gautam 
Type: Summary 
Category: LSC 
Subject: MICH locking tutorial with Gautam 

Today we went through LSC locking mechanics with Gautam and as a "Hello World" example, worked on locking michelson cavity.


MICH settings changed:

  • Gautam at some point added 9 dB attenuation filters in MICH filter module in LSC to match the 9 dB pre-amplifier before digitization.
  • This required changing teh trigger thresholds, C1:LSC-MICH_TRIG_THRESH_ON and C1:LSC-MICH_TRIG_THRESH_OFF.
  • We looked at C1:LSC-AS55_Q_ERR_DQ and C1:LSC-ASDC_OUT_DQ on ndscope.
  • The zero crossings in AS55_Q correspond to ASDC going to zero. We found the threshold values of ASDC by finding the linear region in zero crossing of AS55_Q.
  • We changed the thresold values to UP: -0.3mW and DOWN -0.05mW. The thresholds were also changed in C1LSC_FM_TRIG.
  • We also set FM2,3,6 and 8 to be triggered on threshold.

We characterized the loop OLTF, found the UGF to be 90 Hz and measured the noise at error and control points.

gautam: one aim of this work was to demonstrate that the "Lock Michelson (dark)" script call from the IFOconfigure screen worked - it did, reliably, after the setting changes mentioned above.

ELOG V3.1.3-